# The $\alpha$–regression for compositional data: a unified framework for standard, spatially-lagged, spatial autoregressive and geographically-weighted regression models

### v3

@@ Abstract
Compositional data-vectors of non-negative components summing to unity-frequently arise in scientific applications where covariates influence the relative proportions of components, yet traditional regression approaches ace challenges regarding the unit-sum constraint and zero values. This paper revisits the $\alpha$–regression framework, which uses a flexible power transformation parameterized by $\alpha$ to interpolate between raw data analysis and log-ratio methods, naturally handling zeros without imputation while allowing data-driven transformation selection. We formulate $\alpha$–regression as a non-linear least squares problem, provide efficient estimation via the Levenberg-Marquardt algorithm, and derive marginal effects for interpretation. The framework is extended to spatial settings through two models: the $\alpha$–spatially lagged X regression model, which incorporates spatial spillover effects via spatially lagged covariates with decomposition into direct and indirect effects, the $\alpha$–spatially autoregressive regression model and the geographically weighted $\alpha$–regression, which allows coefficients to vary spatially for capturing local relationships. Applications to two real data sets illustrate the performance of the models and showcase that spatial extensions capture the spatial dependence and improve the predictive performance.
@@ 1 Introduction
Compositional data are characterized as vectors of non-negative components constrained to sum to a constant, conventionally normalized to unity. The sample space of such data is defined by the standard simplex
@@
Compositional data structures arise across diverse scientific domains, as evidenced by the substantial body of methodological literature devoted to their rigorous statistical analysis [^1] For an extensive compilation of domain-specific applications involving compositional data, see (Tsagris and Stewart, 2020)..
The methodological imperative to develop models specifically calibrated for compositional data has catalyzed considerable innovation, particularly in the contemporary statistical literature. The foundational framework was established by Aitchison (2003)—subsequently designated as Aitchison’s model—predicated upon log-ratio transformations, thereby inaugurating the log-ratio approach (LRA). This methodology was subsequently refined by Egozcue et al. (2003), who implemented an isometric log-ratio transformation to preserve geometric properties. In contrast, the stay-in-the-simplex approach employs probability distributions and regression structures intrinsically defined on the simplex manifold. Notably, Dirichlet regression has been extensively utilized within compositional frameworks Gueorguieva et al. (2008), Hijazi and Jernigan (2009), Melo et al. (2009). Furthermore, Iyengar and Dey (2002) investigated the generalized Liouville distribution family, which accommodates negative or heterogeneous correlation structures, thereby extending beyond the restrictive positive correlation constraint of Dirichlet distributions. A less theoretically justified, yet occasionally employed strategy involves disregarding the unit-sum constraint and treating compositional data within a Euclidean framework—an approach designated as raw data analysis (RDA) (Baxter, 2001, Baxter et al., 2005). A fourth methodological paradigm employs the $\alpha$–transformation family (Tsagris et al., 2011), which interpolates continuously between the RDA and LRA, thereby affording enhanced model flexibility while accommodating zero components naturally.
@@
Concerning spatial autocorrelation structures, the spatially lagged X (SLX) model represents a parsimonious specification incorporating spatial dependence exclusively through exogenous covariates, thereby excluding spatial lags of the dependent variable (LeSage and Pace, 2009, Elhorst, 2014). The spatial autoregressive (SAR) model, analogous to temporal autoregressive processes, posits that observations are influenced by proximate spatial neighbors. Specifically, the SAR model expresses the dependent variable as a function of both explanatory covariates and a spatially weighted average of neighboring dependent variable realizations. Nguyen et al. (2021) developed a SAR specification employing the ilr transformation for compositional responses. Geographically weighted regression (GWR) constitutes a local regression methodology designed to capture spatially heterogeneous relationships. In contrast to conventional regression, which assumes parameter stationarity, GWR permits spatial nonstationarity through location-specific coefficient estimation.
The integration of the spatial regression framework within the compositional data analysis represents a relatively narrow research area. Leininger et al. (2013) synthesized GWR with hierarchical Bayesian models for zero-inflated compositional data, incorporating spatial priors to accommodate local variation. Yoshida et al. (2021) implemented the isometric log-ratio transformation prior to GWR estimation, thereby preserving inter-component relationships while facilitating spatially varying parameter estimation. Recently, Clarotto et al. (2022) introduced a novel power transformation, conceptually analogous to the $\alpha$–transformation, specifically calibrated for geostatistical modeling of compositional data.
This paper adopts a pragmatic methodological stance, particularly tailored to regression with compositional data. The principal contribution of this paper is a unified framework for regression modelling of compositional data. The $\alpha$–regression (Tsagris, 2015b), was proposed as a generalization of Aitchison’s log-ratio regression (Aitchison, 2003) that naturally accommodates zero components, while offering flexibility via the $\alpha$–transformation. Based on this, extensions to accomodate spatial correlations are developed.
@@
[…]
@@ 2.1 The $\alpha$–transformation
In effect, $y_{\alpha}$ which resembles a Box–Cox style mapping. The result $y_{\alpha}$ is an unconstrained vector in Euclidean space, suitable for standard multivariate statistical techniques. When $\alpha=1$, the transformation corresponds (up to scaling) to raw data analysis (RDA). When $\alpha=-1$, the transformation is aligned with RDA as well, but using the inverse of the compositional data. As $\alpha\to 0$, the transformation converges to the ilr transformation used in log-ratio analysis (LRA)
@@ 2.2.1 Limiting case of $\alpha\rightarrow 0$
which corresponds to the regression after the centered log-ratio transformation [the ilr transformation (5) without the right multiplication by the Helmert matrix]. This implies that there are $D$ vectors of $\mathbf{\beta}$ regression coefficients. But, since the first set of regression coefficients equals zero, if we subtract this vector from the rest of the $\mathbf{\beta}$ vectors we end up with the regression coefficients of the additive log-ratio (alr) regression
@@ 2.2.2 Choosing $\alpha$
In the regression setting the optimal value of $\alpha$ is data-driven, and there are two ways to estimate its value. The first is to minimize the Kullback-Leibler divergence (KLD) between the observed and fitted compositions $\text{KLD}\left(\mathbf{y},\mathbf{\mu}\right)=\sum_{i=1}^{n}\sum_{j=1}^{D}\mathbf{y}_{ij}\log{y_{i}j/\mu_{ij}}$. This results in a double minimization problem. For a given value of $\alpha$ one must minimize the SSE (7) in order to obtain the regression coefficients and then minimize the KLD with respect to $\alpha$ to obtain the optimal value of $\alpha$. With the choice of the KLD, the value of $\alpha$ is independent of the SSE, since the SSE is not comparable across the different values of $\alpha$.
The second option is to examine $\alpha$ as a hyper-parameter whose value is chosen by minimizing the KLD via CV, e.g. 10-fold CV. (Tsagris, 2015b).
@@ 2.2.3 Marginal effects
To account for the difficult interpretation of the regression coefficients, the marginal effects (MEs), given below, may be used
$\text{ME}_{k}=\frac{\partial\mu_{i}}{\partial x_{k}}=\left\{\begin{array}[]{ll}-\mu_{1}\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}&\text{for}\ \ i=1\\ \mu_{i}\left(\beta_{i-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}\right)&\text{for}\ \ i=2,\ldots,D\end{array}\right\},$ (10)
@@ 2.2.4 Standard error of the MEs
$\text{Var}(g(\mathbf{\theta}))\approx\mathbf{J}\text{Var}(\mathbf{\theta})\mathbf{J}^{\top}$
@@
$m_{i\ell k}=\frac{\partial\mu_{i\ell}}{\partial x_{ik}}$ (11)
The derivatives of $\mu$ with respect to $\mathbf{\beta}$ are:
@@
The Jacobian of the ME for the reference component ($\ell=1$) is
where $\delta_{sk}$ is the Kronecker delta ($\delta_{sk}=1$ if $s=k$, 0 otherwise).
@@ 2.3 Consistency and asymptotic normality
To establish the large-sample properties of the $\alpha$-regression estimator $\hat{\theta}_{n}$, we impose the following assumptions or regularity conditions. The four structural and consistency assumptions delineated below are necesary for the consistency.
@@
The next three assumptions are necessary for asymptotic normality.
@@ 2.4 Robust extensions of the $\alpha$–regression
. The $\alpha$–regression is based upon minimization of the $L_{2}$ norm (7). In a similar fashion one may choose to minimize the sum of the absolute deviations, yielding the $\alpha$–minimum absolute deviations ($\alpha$–MAD) regression
$\text{MAD}\left(\mathbf{Y},\mathbf{X};\alpha,\mathbf{B}\right)=\sum_{i=1}^{n}\|\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\|_{1}=\sum_{i=1}^{n}\sum_{j=1}^{d}|\mathbf{y}_{ij,\alpha}-\mathbf{\mu}_{ij,\alpha}|.$
Following Tsagris (2025b) the $\alpha$–MAD regression was formulated as a univariate regression problem, by using the vectorization operation for the responses and by constructing the design matrix in a suitable manner. To make the estimation efficient the command nlrq() from the R package quantreg (Koenker and others, 2024) was utilized. This approach exhibits dependence on initialization and does not ensure convergence. Alternative models include alternative loss functions. Instead of the $L_{1}$ and $L_{2}$ norms, one may use Tukey’s biweight loss function (Tukey, 1960), Hampel’s loss function (Hampel, 1974) or Barron’s general loss function (Barron, 2019), all available in the R package gslnls (Chau, 2025). A practical limitation of these approaches is the increased computational cost.
@@ 2.5 The $\alpha$–regression with continuous and compositional predictors
$\mu_{i,\alpha^{\prime}}=\begin{cases}\displaystyle\frac{1}{1+\sum_{j=1}^{D}e^{\bm{x}^{\top}\bm{\beta}_{j}+\bm{s}_{\alpha^{\prime}}^{\top}\bm{\gamma}_{j}}}&\text{for }i=1\\[12.91663pt] \displaystyle\frac{e^{\bm{x}^{\top}\bm{\beta}_{i}+\bm{s}_{\alpha^{\prime}}^{\top}\bm{\gamma}_{i}}}{1+\sum_{j=1}^{D}e^{\bm{x}^{\top}\bm{\beta}_{j}+\bm{s}_{\alpha^{\prime}}^{\top}\bm{\gamma}_{j}}}&\text{for }i=2,\ldots,D.\end{cases}$ (15)
@@
$\text{SSE}\left(\bm{Y},\bm{X};\alpha,\alpha^{\prime},\bm{B}\right)=\sum_{i=1}^{n}\|\bm{y}_{i,\alpha}-\bm{\mu}_{i,\alpha^{\prime}}\|_{2}=\sum_{i=1}^{n}\left(\bm{y}_{i,\alpha}-\bm{\mu}_{i,\alpha^{\prime}}\right)^{\top}\left(\bm{y}_{i,\alpha}-\bm{\mu}_{i,\alpha^{\prime}}\right).$ (16)
@@ 3.1.3 Spatial MEs
$DE=\frac{\partial\mu_{i}}{\partial x_{k}}=\begin{cases}-\mu_{1}\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}&\text{for }i=1\\ \mu_{i}\left(\beta_{i-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}\right)&\text{for }i=2,\ldots,D.\end{cases}$ (19)
@@
$IDE=\frac{\partial\mu_{i}}{\partial(\mathbf{W}\mathbf{x})_{k}}=\begin{cases}-\mu_{1}\sum_{j=1}^{d}\gamma_{jk}\mu_{j+1}&\text{for }i=1\\ \mu_{i}\left(\gamma_{i-1,k}-\sum_{j=1}^{d}\gamma_{jk}\mu_{j+1}\right)&\text{for }i=2,\ldots,D.\end{cases}$ (20)
@@
$\frac{\partial\mu_{i}}{\partial x_{k}}+\frac{\partial\mu_{1}}{\partial(\mathbf{W}\mathbf{x})_{k}}=\begin{cases}-\mu_{1}\sum_{j=1}^{d}(\beta_{jk}+\gamma_{jk})\mu_{j+1}&\text{for }i=1\\ \mu_{i}\left[(\beta_{i-1,k}+\gamma_{i-1,k})-\sum_{j=1}^{d}(\beta_{jk}+\gamma_{jk})\mu_{j+1}\right]&\text{for }i=2,\ldots,D.\end{cases}$ (21)

### v4

@@ Abstract
Compositional data–vectors of non-negative components summing to unity–frequently arise in scientific applications where covariates influence the relative proportions of components, yet traditional regression approaches ace challenges regarding the unit-sum constraint and zero values. This paper revisits the $\alpha$–regression framework, which uses a flexible power transformation parameterized by $\alpha$ to interpolate between raw data analysis and log-ratio methods, naturally handling zeros without imputation while allowing data-driven transformation selection. We formulate $\alpha$–regression as a non-linear least squares problem, study its asymptotic properties, provide efficient estimation via the Levenberg-Marquardt algorithm, and derive marginal effects for interpretation. The framework is extended to spatial settings through three models: the $\alpha$–spatially lagged X regression model, which incorporates spatial spillover effects via spatially lagged covariates with decomposition into direct and indirect effects, the $\alpha$–spatially autoregressive regression model and the geographically weighted $\alpha$–regression, which allows coefficients to vary spatially for capturing local relationships. Applications to two real data sets illustrate the performance of the models and showcase that spatial extensions capture the spatial dependence and improve the predictive performance.
@@ 1 Introduction
Compositional data are characterized as vectors of non-negative components constrained to sum to a constant, conventionally normalized to unity. Compositional data structures arise across diverse scientific domains, as evidenced by the substantial body of methodological literature devoted to their rigorous statistical analysis [^1] For an extensive compilation of domain-specific applications involving compositional data, see (Tsagris and Stewart, 2020).. The sample space of such data is defined by the standard simplex
@@
The methodological imperative to develop models specifically calibrated for compositional data has catalyzed considerable innovation, particularly in the contemporary statistical literature. The foundational framework was established by Aitchison (2003)–subsequently designated as Aitchison’s model–predicated upon log-ratio transformations, thereby inaugurating the log-ratio analysis (LRA). This methodology was subsequently refined by Egozcue et al. (2003), who implemented an isometric log-ratio (ilr) transformation to preserve geometric properties. In contrast, the stay-in-the-simplex approach employs probability distributions and regression structures intrinsically defined on the simplex manifold. Notably, Dirichlet regression has been extensively utilized within compositional frameworks Gueorguieva et al. (2008), Hijazi and Jernigan (2009), Melo et al. (2009). Furthermore, Iyengar and Dey (2002) investigated the generalized Liouville distribution family, which accommodates negative or heterogeneous correlation structures, thereby extending beyond the restrictive positive correlation constraint of Dirichlet distributions. A less theoretically justified, yet occasionally employed strategy involves disregarding the unit-sum constraint and treating compositional data within a Euclidean framework—an approach designated as raw data analysis (RDA) (Baxter, 2001, Baxter et al., 2005). A fourth methodological paradigm employs the $\alpha$–transformation family (Tsagris et al., 2011), which interpolates continuously between the RDA and LRA, thereby affording enhanced model flexibility while accommodating zero components naturally.
@@
Concerning spatial autocorrelation structures, the spatially lagged X (SLX) model represents a parsimonious specification incorporating spatial dependence exclusively through exogenous covariates, thereby excluding spatial lags of the dependent variable (LeSage and Pace, 2009, Elhorst, 2014). The spatial autoregressive (SAR) model (Kazar and Celik, 2012, Shi et al., 2025), analogous to temporal autoregressive processes, posits that observations are influenced by proximate spatial neighbors. Specifically, the SAR model expresses the dependent variable as a function of both explanatory covariates and a spatially weighted average of neighboring dependent variable realizations. Geographically weighted regression (GWR) constitutes a local regression methodology designed to capture spatially heterogeneous relationships (Brunsdon et al., 1996). In contrast to conventional regression, which assumes parameter stationarity, GWR permits spatial nonstationarity through location-specific coefficient estimation.
The integration of the spatial regression framework within the compositional data analysis represents a relatively narrow research area [^2] The literature in spatial modelling contains more works, but most of them rely on log-ratio transformations prior to the application of standard spatial models.. Leininger et al. (2013) synthesized hierarchical Bayesian models for zero-inflated compositional data, incorporating spatial random effects to accommodate local variation. Nguyen et al. (2021) developed a SAR specification employing the ilr transformation for compositional responses. Yoshida et al. (2021) implemented the ilr transformation prior to GWR estimation, thereby preserving inter-component relationships while facilitating spatially varying parameter estimation. Clarotto et al. (2022) introduced a novel power transformation, conceptually analogous to the $\alpha$–transformation, specifically calibrated for geostatistical modeling of compositional data.
This paper adopts a pragmatic methodological stance, particularly tailored to regression with compositional data. The principal contribution of this paper is a unified framework for regression modelling of compositional data. The $\alpha$–regression (Tsagris, 2015b), was proposed as a generalization of Aitchison’s log-ratio regression (Aitchison, 2003) that naturally accommodates zero components, while offering flexibility via the $\alpha$–transformation. At first we review the $\alpha$–regression model, examine it as a non-linear least squares minimization problem and use a modified Levenberg-Marquardt algorithm, that is comutationally efficient, to estimate the regression coefficients. We provide formulas for the marginal effects (MEs) of the covariates, including their asymptotic variance, We then establish the consistency and the asymptotic normality of the regression coefficients. Closing with the simple $\alpha$–regression, we discuss robust extensions and a simple method to incorporate compositional and Euclidean predictors.
@@
The next section discusses the $\alpha$–regression, while section 3 generalizes this model to its spatial regression versions. Section 4 illustrates the performance of all regression models on two real datasets and Section 5 concludes the paper.
@@ 2.1 The $\alpha$–transformation
In effect, $y_{\alpha}$ which resembles a Box–Cox style mapping. The resulting $y_{\alpha}$ is an unconstrained vector in Euclidean space, suitable for standard multivariate statistical techniques. When $\alpha=1$, the transformation corresponds (up to scaling) to raw data analysis (RDA). When $\alpha=-1$, the transformation is aligned with RDA as well, but using the inverse of the compositional data. As $\alpha\to 0$, the transformation converges to the ilr transformation used in LRA.
@@ 2.2.1 Limiting case of $\alpha\rightarrow 0$
which corresponds to the regression after the centered log-ratio transformation [the ilr transformation (5) without the right multiplication by the Helmert matrix]. This implies that there are $D$ vectors of $\mathbf{\beta}$ regression coefficients. But, since the set of regression coefficients sums to zero, if we subtract the first coefficient from the rest of the $\mathbf{\beta}_{s}$ we end up with the regression coefficients of the additive log-ratio (alr) regression
@@ 2.2.2 Choosing $\alpha$
In the regression setting the optimal value of $\alpha$ is data-driven, and there are two ways to estimate its value. The first is to minimize the Kullback-Leibler divergence (KLD) between the observed and fitted compositions $\text{KLD}\left(\mathbf{y},\mathbf{\mu}\right)=\sum_{i=1}^{n}\sum_{j=1}^{D}\mathbf{y}_{ij}\log{y_{ij}/\mu_{ij}}$. This results in a double minimization problem. For a given value of $\alpha$ one must minimize the SSE (7) in order to obtain the regression coefficients and then minimize the KLD with respect to $\alpha$ to obtain the optimal value of $\alpha$. With the choice of the KLD, the value of $\alpha$ is independent of the SSE, since the SSE is not comparable across the different values of $\alpha$. The second option is to examine $\alpha$ as a hyper-parameter whose value is chosen by minimizing the KLD via CV, e.g. 10-fold CV. (Tsagris, 2015b).
@@ 2.2.3 Marginal effects
To account for the difficult interpretation of the regression coefficients, the MEs, given below, may be used
$\text{ME}_{ik}=\frac{\partial\mu_{i}}{\partial x_{k}}=\left\{\begin{array}[]{ll}-\mu_{1}\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}&\text{for}\ \ i=1\\ \mu_{i}\left(\beta_{i-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}\right)&\text{for}\ \ i=2,\ldots,D\end{array}\right\},$ (10)
@@ 2.2.4 Standard error of the MEs
$\text{Var}(g(\mathbf{\theta}))\approx\mathbf{J}\text{Var}(\mathbf{\theta})\mathbf{J}^{\top},$
@@
$ME_{i\ell k}=m_{i\ell k}=\frac{\partial\mu_{i\ell}}{\partial x_{ik}}.$ (11)
The Jacobian of the ME for the reference component ($\ell=1$) is
where $\delta_{sk}$ is the Kronecker delta ($\delta_{sk}=1$ if $s=k$, 0 otherwise), and the derivatives of $\mu$ with respect to $\mathbf{\beta}$ are:
@@ 2.3 Consistency and asymptotic normality
To establish the large-sample properties of the $\alpha$-regression estimator $\hat{\theta}_{n}$, we impose the following assumptions or regularity conditions. The four structural assumptions delineated below are necesary to establish the consistency of the regression coefficients.
@@
The next three assumptions are necessary to establish the asymptotic normality of the regression coefficients.
@@ 2.4 Robust extensions of the $\alpha$–regression
The $\alpha$–regression is based upon minimization of the $L_{2}$ norm (7). In a similar fashion one may choose to minimize the sum of the absolute deviations, yielding the $\alpha$–minimum absolute deviations ($\alpha$–MAD) regression
$\text{MAD}\left(\mathbf{Y},\mathbf{X};\alpha,\mathbf{B}\right)=\sum_{i=1}^{n}\sum_{j=1}^{d}|\mathbf{y}_{ij,\alpha}-\mathbf{\mu}_{ij,\alpha}|.$
Following Tsagris (2025b) the $\alpha$–MAD regression was formulated as a univariate regression problem, by using the vectorization operation for the responses and by constructing the design matrix in a suitable manner. To make the estimation efficient the command nlrq() from the R package quantreg (Koenker and others, 2024) was utilized. This approach exhibits dependence on initialization and does not ensure convergence. Alternative models include alternative loss functions. Instead of the $L_{2}$ norm, one may use the $L_{1}$ norm $\sum_{i=1}^{n}\|\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\|_{1}$ leading to the $\alpha$–spatial median regression [^5] One option for this regression is to use R’s built-in optimizers or iteratively reweighted least squares. The second option is computationally expensive as well, given the complexity of the derivatives involved.. Tukey’s biweight loss function (Tukey, 1960), Hampel’s loss function (Hampel, 1974) or Barron’s general loss function (Barron, 2019), all available in the R package gslnls (Chau, 2025). A practical limitation of these approaches is the increased computational cost.
@@ 2.5 The $\alpha$–regression with continuous and compositional predictors
$\mu_{i}^{\alpha^{\prime}}=\begin{cases}\displaystyle\frac{1}{1+\sum_{j=1}^{D}e^{\bm{x}^{\top}\bm{\beta}_{j}+\bm{s}_{\alpha^{\prime}}^{\top}\bm{\gamma}_{j}}}&\text{for }i=1\\[12.91663pt] \displaystyle\frac{e^{\bm{x}^{\top}\bm{\beta}_{i}+\bm{s}_{\alpha^{\prime}}^{\top}\bm{\gamma}_{i}}}{1+\sum_{j=1}^{D}e^{\bm{x}^{\top}\bm{\beta}_{j}+\bm{s}_{\alpha^{\prime}}^{\top}\bm{\gamma}_{j}}}&\text{for }i=2,\ldots,D.\end{cases}$ (15)
@@
$\text{SSE}\left(\bm{Y},\bm{X};\alpha,\alpha^{\prime},\bm{B}\right)=\sum_{i=1}^{n}\|\bm{y}_{i,\alpha}-\bm{\mu}_{i,\alpha^{\prime}}\|_{2}=\sum_{i=1}^{n}\left(\bm{y}_{i,\alpha}-\bm{\mu}_{i,\alpha}^{\alpha^{\prime}}\right)^{\top}\left(\bm{y}_{i,\alpha}-\bm{\mu}_{i,\alpha}^{\alpha^{\prime}}\right).$ (16)
@@ 3.1.3 Spatial MEs
$DSME_{ik}=\frac{\partial\mu_{i}}{\partial x_{k}}=\begin{cases}-\mu_{1}\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}&\text{for }i=1\\ \mu_{i}\left(\beta_{i-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}\right)&\text{for }i=2,\ldots,D.\end{cases}$
@@
$IDSME_{ik}=\frac{\partial\mu_{i}}{\partial(\mathbf{W}\mathbf{x})_{k}}=\begin{cases}-\mu_{1}\sum_{j=1}^{d}\gamma_{jk}\mu_{j+1}&\text{for }i=1\\ \mu_{i}\left(\gamma_{i-1,k}-\sum_{j=1}^{d}\gamma_{jk}\mu_{j+1}\right)&\text{for }i=2,\ldots,D.\end{cases}$
@@
$TSME_{ik}=\frac{\partial\mu_{i}}{\partial x_{k}}+\frac{\partial\mu_{1}}{\partial(\mathbf{W}\mathbf{x})_{k}}=\begin{cases}-\mu_{1}\sum_{j=1}^{d}(\beta_{jk}+\gamma_{jk})\mu_{j+1}&\text{for }i=1\\ \mu_{i}\left[(\beta_{i-1,k}+\gamma_{i-1,k})-\sum_{j=1}^{d}(\beta_{jk}+\gamma_{jk})\mu_{j+1}\right]&\text{for }i=2,\ldots,D.\end{cases}$